Question 2


A perfectly competitive firm faces the short-run cost schedule shown in Table 1.


Output


Total Cost


12


1


20


2


26


3


32


Table 1


4


40


567


68


93


a) Calculate average total cost (ATC=TC/Q), marginal cost (MC=ATC/AQ) and marginal revenue (MR-ATR/AQ) for each level of output. The price per unit of output is £16.


[5 marks]


b) Plot ATC, MC and MR on a graph and mark the profit-maximising output. At what output is profit maximised?


[5 marks]


8


122


BUS108 (2021)


c) How much profit/loss is made at the optimum level of output?


Page 3


[5 marks]


d) Assume market price declines to £9 per unit. If the firm's average variable cost is £9.5, should the firm shut down in the short run? In the long run? Explain.


[5 marks]


e) If the firm is typical of other firms, what price will it charge in the long run? Explain.


[5 marks]


1
Expert's answer
2021-05-16T17:46:04-0400

A perfectly competitive firm maximizes profit by prducing output at level where P = MC

Total Revenue = Price * Quantity

Marginal Revenue = Change in Total Revenue / Change in Quantity

Marginal Cost = Change in Total Cost / Change in Quantity

(a)




(b)



At output of 32 units P or MR > MC as 16 > 4.83


(c)

At 32 units of output

TR = 512

TC = 122

Maximum Profit "= 512 - 122 = 390"


(d)For a firm to reach a point of shutdown, the condition is that the variable cost(average) must be equal/less than the price(revenue in average).

With the given data, the AR is at £9, which is the price. But the variable cost(average) is higher than this at £9.5. Thus, there is no need for a shutdown, at this point. 

In the long run, if the variable cost(average) goes below £9.5 or the price does not rise, then a need for requirement might occur. But it can be avoided if the firm starts charging a price that is higher. 


(e)The price in the long will be equal to the total cost(average) and also with the marginal cost. 

Thus, in this case, it will be 

"\frac{Total\space of\space costs}{total\space of\space output}""=\frac{902}{92}=£9.39"

Thus, a price of "\u00a39.39" will be charged.